OpenStudy (anonymous):

Solve the following diffusion problem. Ut=4Uxx U(0,t)=0 U(3,t)=0 U(x,0)=sin(2pix/3)-2sin(pix)+7sin(5pix/3)

OpenStudy (anonymous):

\[u _{t}=4u _{xx}\]

OpenStudy (anonymous):

\[u(0,t)=0 , u(3,t)=0\]

OpenStudy (anonymous):

\[u(x,0)=\sin(\frac{ 2 \pi x }{ 3 })-2\sin(\pi x)+7\sin(\frac{ 5 \pi x }{ 3 })\]